### v2

# An extension of Viennot’s shadow to rook placements via orbit harmonics

Abstract.
For fixed positive integers $n,m,r$ with $r\leq\min(n,m)$, let ${\mathrm{Mat}}_{n\times m}({\mathbb{C}})$ be the affine space of $n\times m$ complex matrices with coordinate ring ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]$. We define a homogeneous ideal $I_{n,m,r}$, where the graded quotient ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ carries an action of the group ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$ by independent permutations of rows and columns. This quotient ring is obtained by applying the orbit harmonics method to matrix loci corresponding to all rook placements of size at least $r$. We show that the quotient ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ admits a standard monomial basis, which is determined by extending rook placements to an element in ${\mathfrak{S}}_{n+m-r}$ and applying Viennot’s shadow line avatar of the Schensted correspondence. We also determine the graded ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-module structure of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$.
1. Introduction
Let ${\mathbf{x}}_{N}=(x_{1},x_{2},\dots,x_{N})$ be a collection of $N$ variables, and let ${\mathbb{C}}[{\mathbf{x}}_{N}]:={\mathbb{C}}[x_{1},x_{2},\dots,x_{N}]$ be the polynomial ring over these variables. For a finite locus of points ${\mathcal{Z}}\subseteq{\mathbb{C}}^{N}$, its vanishing ideal is
(1.1) ${\mathbf{I}}({\mathcal{Z}}):=\{f\in{\mathbb{C}}[{\mathbf{x}}_{N}]:f(z)=0\text{ for all }z\in{\mathcal{Z}}\}.$
The coordinate ring ${\mathbb{C}}[{\mathcal{Z}}]$ can then be identified with ${\mathbb{C}}[{\mathbf{x}}_{N}]/{\mathbf{I}}({\mathcal{Z}})$. The method of orbit harmonics replaces the ideal ${\mathbf{I}}({\mathcal{Z}})$ with its associated graded ideal ${\mathrm{gr}}\,{\mathbf{I}}({\mathcal{Z}})$, extending the identification of (ungraded) vector spaces to
(1.2) ${\mathbb{C}}[{\mathcal{Z}}]\cong{\mathbb{C}}[{\mathbf{x}}_{N}]/{\mathbf{I}}({\mathcal{Z}})\cong{\mathbb{C}}[{\mathbf{x}}_{N}]/{\mathrm{gr}}\,{\mathbf{I}}({\mathcal{Z}})=:R({\mathcal{Z}}).$
If $\mathcal{Z}$} is invariant under the action of a finite matrix group $G\subseteq\mathrm{GL}_{N}({\mathbb{C}})$, this isomorphism can be treated as an isomorphism of $G$-modules, and $R({\mathcal{Z}})$ has the additional structure as a graded $G$-module.
Geometrically, the method of orbit harmonics linearly deforms the locus $\mathcal{Z}$} to a scheme of multiplicity $|{\mathcal{Z}}|$ supported at the origin. In the picture below, this deformation is drawn schematically for $\mathcal{Z}$}, which is a locus of six points stable under the action of ${\mathfrak{S}}_{3}$ generated by reflections across the three lines.
$\bullet\bullet\bullet\bullet\bullet\bullet\bullet$
The method of orbit harmonics dates back to Kostant  [6]. It has seen its application in numerous settings, including cohomology theory  [3], Macdonald theory  [4, 5], cyclic sieving [10], Donaldson–Thomas invariants  [11], and Ehrhart theory  [12]. In cases where the locus $\mathcal{Z}$} has nice combinatorial structures or symmetries, one can usually expect that the algebraic properties of $R({\mathcal{Z}})$ are governed by the combinatorial properties of $\mathcal{Z}$}.
Let ${\mathbf{x}}_{n\times m}=(x_{i,j})_{1\leq i\leq n,\,1\leq j\leq m}$ be an $n\times m$ matrix of variables. Let ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]$ be the polynomial ring over these variables and ${\mathrm{Mat}}_{n\times m}({\mathbb{C}})$ be the affine space of $n\times m$ complex matrices. The application of orbit harmonics to finite matrix loci ${\mathcal{Z}}\subseteq{\mathrm{Mat}}_{n\times m}({\mathbb{C}})$ was initiated by Rhoades [13]. He considered the case $n=m$ with the point locus ${\mathcal{Z}}={\mathfrak{S}}_{n}\subseteq{\mathrm{Mat}}_{n\times n}({\mathbb{C}})$ of $n\times n$ permutation matrices. This permutation matrix locus carries an action of ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{n}$ by left and right multiplication. Algebraic properties of $R({\mathfrak{S}}_{n})$ are governed by the longest increasing subsequences in ${\mathfrak{S}}_{n}$ and Viennot’s shadow line construction. Liu [7] extended this work to the locus ${\mathcal{Z}}={\mathbb{Z}}_{r}\wr{\mathfrak{S}}_{n}$ of $r$-colored permutation matrices. Liu, Ma, Rhoades, Zhu  [8] further studied the matrix loci ${\mathcal{Z}}=\{w\in{\mathfrak{S}}_{n}:w^{2}=1\}\subseteq{\mathrm{Mat}}_{n\times n}({\mathbb{C}})$ of involutions and the matrix loci ${\mathcal{Z}}=\mathcal{PM}_{n}=\{w\in{\mathcal{M}}_{n}:w(i)\neq i\text{ for all }i\}$ of fixed-point-free involutions.
In this paper, we consider a matrix locus formed by rook placements. For positive integers $n$ and $m$, a rook placement $\mathcal{R}$} on the $n\times m$ board is a subset of $n]\times[m]$ such that $(i,j)\in{\mathcal{R}}$ and $(i^{\prime},j^{\prime})\in{\mathcal{R}}$ implies $i\neq i^{\prime}$ and $j\neq j^{\prime}$. We call each element $(i,j)\in{\mathcal{R}}$ a rook of $\mathcal{R}$}. Define the size $|{\mathcal{R}}|$ of $\mathcal{R}$} by
$|{\mathcal{R}}|\coloneqq\text{the number of rooks of ${\mathcal{R}}$}.$
Identify each rook placement $\mathcal{R}$} on the $n\times m$ board with the $n\times m$ matrix $(a_{i,j})_{1\leq i\leq n,\,1\leq j\leq m}$ given by
$a_{i,j}=\begin{cases}1,&\text{if $(i,j)\in{\mathcal{R}}$}\\ 0,&\text{otherwise.}\end{cases}$
Write ${\mathcal{Z}}_{n,m,r}\subseteq{\mathrm{Mat}}_{n\times m}({\mathbb{C}})$ for the set of all rook placements on the $n\times m$ board of size $r$, where $r\leq\min\{m,n\}$. We are interested in the finite locus of all rook placements with at least $r$ rooks:
${\mathcal{UZ}}_{n,m,r}\coloneqq\bigsqcup_{r^{\prime}=r}^{\min\{m,n\}}{\mathcal{Z}}_{n,m,r^{\prime}}\subseteq{\mathrm{Mat}}_{n\times m}({\mathbb{C}}).$
The locus ${\mathcal{UZ}}_{n,m,r}$ is closed under the action of ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$ by left and right multiplication, and the quotient $R({\mathcal{UZ}}_{n,m,r})$ is a graded ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-module. Our results on this quotient include:
  • We give an explicit generating set of ${\mathrm{gr}}\,{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$. See Definition 3.1 and Theorem 4.12.
  • We show (Theorem 4.12) that $R({\mathcal{UZ}}_{n,m,r})$ admits a monomial basis ${\mathfrak{es}}({\mathcal{R}})$ indexed by rook placements in ${\mathcal{UZ}}_{n,m,r}$ (see Definition 4.8), which is defined by extending each rook placement to a permutation in ${\mathfrak{S}}_{n+m-r}$ and applying Viennot’s shadow construction. We further show that this is the standard monomial basis with respect to diagonal monomial orders (Definition 4.7).
  • We describe the graded ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-module structure of $R({\mathcal{UZ}}_{n,m,r})$.
The rest of the paper is structured as follows. In Section 2 we give background material on Gröbner theory, orbit harmonics, and representations of ${\mathfrak{S}}_{n}$. In Section 3 we define the ideal $I_{n,m,r}$, and study its algebraic structure. In Section 4 we define Viennot’s shadow line construction on ${\mathfrak{S}}_{n}$ and our extension of this construction to ${\mathcal{UZ}}_{n,m,r}$. We show that this construction gives the standard monomial basis of $R({\mathcal{UZ}}_{n,m,r})$ and prove that the ideal we define in Section 3 is exactly ${\mathrm{gr}}\,{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$. In Section 5 we discuss the structure of the graded ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-module $R({\mathcal{UZ}}_{n,m,r})$. We close in Section 6 with conjectures and possible directions of future research.
2. Background
2.1. Gröbner Theory
Let ${\mathbf{x}}_{N}=(x_{1},x_{2},\dots x_{N})$ be a sequence of $N$ variables, and let ${\mathbb{C}}[{\mathbf{x}}_{N}]$ be the polynomial ring over these variables. A total order << on the set of monomials in ${\mathbb{C}}[{\mathbf{x}}_{N}]$ is called a monomial order if
  • $1\leq m$ for every monomial $m\in{\mathbb{C}}[{\mathbf{x}}_{N}]$
  • for monomials $m,\,m_{1},\,m_{2}$, we have that $m_{1}<m_{2}$ implies $mm_{1}<mm_{2}$.
Let << be a monomial order. For any nonzero polynomial $f\in{\mathbb{C}}[{\mathbf{x}}_{N}]$, the initial monomial $\text{in}_{<}f$ of $f$ is the largest monomial in $f$ with respect to <<. Given an ideal $I\subseteq{\mathbb{C}}[{\mathbf{x}}_{N}]$, its initial ideal is defined to be
(2.1) $\text{in}_{<}I=\langle\text{in}_{<}f:f\in I,f\neq 0\rangle.$
A monomial $m\in{\mathbb{C}}[{\mathbf{x}}_{N}]$ is called a standard monomial of $I$ if it is not an element of $\text{in}_{<}I$. It is well known that
(2.2) $\{m+I:m\text{ is a standard monomial of }I\}$
forms a basis for the vector space ${\mathbb{C}}[{\mathbf{x}}_{N}]/I$. This is called the standard monomial basis. See [2] for further details on Gröbner theory.
2.2. Orbit Harmonics
Let ${\mathcal{Z}}\subset{\mathbb{C}}^{N}$ be a finite locus of points in the affine $N$-space. The vanishing ideal ${\mathbf{I}}({\mathcal{Z}})\subseteq{\mathbb{C}}[{\mathbf{x}}_{N}]$ is the ideal
(2.3) ${\mathbf{I}}({\mathcal{Z}})=\{f\in{\mathbb{C}}[{\mathbf{x}}_{N}]:f(z)=0\text{ for all }z\in{\mathcal{Z}}\}.$
Identifying ${\mathbb{C}}[{\mathcal{Z}}]$ with the vector space of all functions from $\mathcal{Z}$} to $\mathbb{C}$}, we obtain
(2.4) ${\mathbb{C}}[{\mathcal{Z}}]\cong{\mathbb{C}}[{\mathbf{x}}_{N}]/{\mathbf{I}}({\mathcal{Z}}).$
Given a nonzero polynomial $f\in{\mathbb{C}}[{\mathbf{x}}_{N}]$, write $f=f_{0}+f_{1}+\dots+f_{d}$, where $f_{i}$ is homogeneous of degree $i$ and $f_{d}\neq 0$. Denote by $\tau(f)$ the top degree homogeneous part of $f$, that is, $\tau(f)=f_{d}$. For any ideal $I\subseteq{\mathbb{C}}[{\mathbf{x}}_{N}]$, its associated graded ideal is
(2.5) ${\mathrm{gr}}\,I=\langle\tau(f):f\in I\rangle.$
Remark 2.1.
Given a set of generators $I=\langle f_{1},f_{2},\dots,f_{r}\rangle$, we have the containment of ideals $\langle\tau(f_{1}),\tau(f_{2}),\dots,\tau(f_{r})\rangle\subseteq{\mathrm{gr}}\,I$. This containment is strict in general. On the other hand, if $\{f_{1},f_{2},\dots,f_{r}\}$ forms a Gröbner basis of $I$ with respect to some graded monomial order, we have the equality $\langle\tau(f_{1}),\tau(f_{2}),\dots,\tau(f_{r})\rangle={\mathrm{gr}}\,I$.
The isomorphism of ungraded vector spaces in Equation (2.4) can be extended to
(2.6) ${\mathbb{C}}[{\mathcal{Z}}]\cong{\mathbb{C}}[{\mathbf{x}}_{N}]/{\mathbf{I}}({\mathcal{Z}})\cong{\mathbb{C}}[{\mathbf{x}}_{N}]/{\mathrm{gr}}\,{\mathbf{I}}({\mathcal{Z}})=:R({\mathcal{Z}}).$
Since the generators of ${\mathrm{gr}}\,{\mathbf{I}}({\mathcal{Z}})$ are all homogeneous, $R({\mathcal{Z}})$ has the additional structure of a graded vector space. Furthermore, if $\mathcal{Z}$} is stable under the action of a matrix group $G\subseteq GL_{N}({\mathbb{C}})$, Equation  (2.6) is also an isomorphism of ungraded $G$-modules, and $R({\mathcal{Z}})$ has the additional structure of a graded $G$-module.
2.3. Schensted Correspondence
Let $n$ be any positive integer. A partition $\lambda$ of $n$ is a sequence of non-increasing positive integers $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{k})$ such that $\sum_{i=1}^{k}\lambda_{i}=n$. We write $\lambda\vdash n$ to indicate that $\lambda$ is a partition of $n$.
For a partition $\lambda\vdash n$, its Young diagram is a diagram with $n$ left-justified boxes, with $\lambda_{i}$ boxes on the $i$-th row. A standard Young Tableau of shape $\lambda$ is a filling of $[n]$ into the boxes in the Young diagram of $\lambda$, so that the numbers increase across each row and down each column. Below on the left is the Young diagram of the partition $\lambda=(4,2,1)\vdash 7$ and on the right is an example of a standard Young tableau of shape $\lambda$.
11 77 22 44 66
Denote by $\mathrm{SYT}(\lambda)$ the set of standard Young tableaux of shape $\lambda$. The Schensted correspondence [15] is a bijection
(2.7) ${\mathfrak{S}}_{n}\xrightarrow{\quad\sim\quad}\bigsqcup_{\lambda\vdash n}\{(P,Q):P,Q\in\mathrm{SYT}(\lambda)\}$
that takes a permutation in ${\mathfrak{S}}_{n}$ to a pair of standard Young tableaux of the same shape. This bijection is usually achieved through an insertion algorithm, details of which can be found in  [14]. In Section 4.1, we will introduce an alternate description of the Schensted correspondence discovered by Viennot [16].
2.4. Representation Theory
[…]
The degree-$n$ component $\Lambda_{n}$ of $\Lambda$ has bases indexed by partitions $\lambda\vdash n$. For example, the complete homogeneous basis of $\Lambda_{n}$ is $\{h_{\lambda}\,:\,\lambda\vdash n\}$ where $h_{\lambda}=h_{\lambda_{1}}h_{\lambda_{2}}\dots$ for $\lambda=(\lambda_{1}\geq\lambda_{2}\geq\dots)$. We will also need the Schur basis $\{s_{\lambda}\,:\,\lambda\vdash n\}$ (see [9] for its definition). If $\lambda=(n)$ one has $s_{(n)}=h_{n}$.
The algebra of doubly symmetric functions is the tensor product $\Lambda\otimes_{{\mathbb{C}}(q)}\Lambda$, which is naturally doubly graded. The Schur basis of the degree $(n,m)$ piece is $\{s_{\lambda}\otimes s_{\mu}\,:\,\lambda\vdash n\,,\,\mu\vdash m\}$.
A symmetric function $F\in\Lambda$ is Schur-positive if it can be written as $F=\sum_{\lambda}c_{\lambda}(q)\cdot s_{\lambda}$ with coefficients $c_{\lambda}(q)\in\mathbb{R}_{\geq 0}[q]$. Similarly, a doubly symmetric function $F^{\prime}\in\Lambda\otimes_{{\mathbb{C}}(q)}\Lambda$ is Schur-positive if it can be written as $F^{\prime}=\sum_{\lambda,\mu}c_{\lambda,\mu}(q)\cdot s_{\lambda}\otimes s_{\mu}$ with coefficients $c_{\lambda,\mu}\in\mathbb{R}_{\geq 0}[q]$. For $F,G\in\Lambda$ (resp. $F,G\in\Lambda\otimes_{{\mathbb{C}}(q)}\Lambda$), we write $F\leq G$ if and only if GFG-F is Schur-positive.
We introduce the truncation operator $\{-\}_{P}$ where $P$ is a condition. Given $F=\sum_{\lambda}c_{\lambda}s_{\lambda}\in\Lambda$ with $c_{\lambda}\in{\mathbb{C}}(q)$, we define $\{F\}_{P}$ by
$\{F\}_{P}\coloneqq\sum_{\text{$\lambda$ satisfies $P$}}c_{\lambda}s_{\lambda}.$
Given $G=\sum_{\lambda,\mu}c_{\lambda,\mu}s_{\lambda}\otimes s_{\mu}\in\Lambda\otimes_{{\mathbb{C}}(q)}\Lambda$ with $c_{\lambda,\mu}\in{\mathbb{C}}(q)$ for all partitions $\lambda,\mu$, we similarly define $\{G\}_{P}$ by
$\{G\}_{P}\coloneqq\sum_{\text{$\lambda,\mu$ both satisfy $P$}}c_{\lambda,\mu}s_{\lambda}\otimes s_{\mu}.$
For instance, we have
$\{G\}_{\lambda_{1}\leq L}=\sum_{\begin{subarray}{c}\lambda_{1}\leq L\\ \mu_{1}\leq L\end{subarray}}c_{\lambda,\mu}s_{\lambda}\otimes s_{\mu}.$
Irreducible representations of ${\mathfrak{S}}_{n}$ are indexed by partitions $\lambda\vdash n$. We write $V^{\lambda}$ for the irreducible ${\mathfrak{S}}_{n}$-module indexed by $\lambda$. Given an ${\mathfrak{S}}_{n}$-module $V$, there exists a unique decomposition $V\cong\bigoplus_{\lambda\vdash n}c_{\lambda}V^{\lambda}$ with $c_{\lambda}\geq 0$. The Frobenius image of $V$ is defined to be
(2.8) ${\mathrm{Frob}}(V)=\sum_{\lambda\vdash n}c_{\lambda}s_{\lambda},$
obtained by replacing irreducible representation $V^{\lambda}$ with the Schur polynomial $s_{\lambda}$. If $V=\bigoplus_{d\geq 0}V_{d}$ is a graded ${\mathfrak{S}}_{n}$-module, the graded Frobenius image of $V$ is
(2.9) ${\mathrm{grFrob}}(V,q)=\sum_{d\geq 0}{\mathrm{Frob}}(V_{d})\cdot q^{d}.$
For integers $n,m\geq 0$, irreducible representations of ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$ are of the form $V^{\lambda}\otimes V^{\mu}$, where $\lambda\vdash n$ and $\mu\vdash m$. Each ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-module $V$ has a unique decomposition $V\cong\bigoplus c_{\lambda,\mu}V^{\lambda}\otimes V^{\mu}$ with $c_{\lambda,\mu}\geq 0$, and we extend the definition of Frobenius images to ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-modules by letting
(2.10) ${\mathrm{Frob}}(V)=\sum_{\begin{subarray}{c}\lambda\vdash n\\ \mu\vdash m\end{subarray}}c_{\lambda,\mu}s_{\lambda}\otimes s_{\mu},$
and define the graded Frobenius images for graded ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$ analogously.
We need the following result about symmetrizers acting on irreducible representations of ${\mathfrak{S}}_{n}$. Let $j<n$ be a positive integer and embed ${\mathfrak{S}}_{j}\subseteq{\mathfrak{S}}_{n}$ by letting it act on the first $j$ letters. This induces an embedding of group algebras ${\mathbb{C}}[{\mathfrak{S}}_{j}]\subseteq{\mathbb{C}}[{\mathfrak{S}}_{n}]$. Let $\eta_{j}\in{\mathbb{C}}[{\mathfrak{S}}_{j}]\subseteq{\mathbb{C}}[{\mathfrak{S}}_{n}]$ be the element that symmetrizes over ${\mathfrak{S}}_{j}$, i.e.
(2.11) $\eta_{j}=\sum_{w\in{\mathfrak{S}}_{j}}w.$
If $V$ is an ${\mathfrak{S}}_{n}$-module, $\eta_{j}$ acts as an operator on $V$, and we have the following lemma which characterizes when $\eta_{j}$ annihilates irreducible ${\mathfrak{S}}_{n}$-modules.
Lemma 2.2.
For $j\leq n$ and $\lambda\vdash n$, we have that $\eta_{j}\cdot V^{\lambda}\neq 0$ if and only if $\lambda_{1}\geq j$.
Proof.
For any irreducible ${\mathfrak{S}}_{n}$-module $V^{\lambda}$, we have
(2.12) $\eta_{j}\cdot V^{\lambda}=(V^{\lambda})^{{\mathfrak{S}}_{j}}=\{v\in V^{\lambda}\,:\,w\cdot v=v\text{ for all }w\in{\mathfrak{S}}_{j}\}.$
Thus, $\eta_{j}\cdot V^{\lambda}\neq 0$ if and only if $(\mathrm{Res}^{{\mathfrak{S}}_{n}}_{{\mathfrak{S}}_{j}}\,V^{\lambda})^{{\mathfrak{S}}_{j}}\neq 0$. This is true if and only if the trivial representation $V^{(j)}$ appears with positive multiplicity in the decomposition of $\mathrm{Res}^{{\mathfrak{S}}_{n}}_{{\mathfrak{S}}_{j}}\,V^{\lambda}$. The branching rule of ${\mathfrak{S}}_{n}$-modules (see [9, 14]) states that
(2.13) $\mathrm{Res}^{{\mathfrak{S}}_{n}}_{{\mathfrak{S}}_{j}}\,V^{\lambda}=\bigoplus_{\begin{subarray}{c}\mu\vdash j,\\ \mu_{i}\leq\lambda_{i}\text{ for all }i\end{subarray}}V^{\mu},$
which implies that $V^{(j)}$ has positive multiplicity in $\mathrm{Res}^{{\mathfrak{S}}_{n}}_{{\mathfrak{S}}_{j}}\,V^{\lambda}$ if and only if $\lambda_{1}\geq j$, completing the proof. ∎
3. The ideal $I_{n,m,r}$ and its structure
Throughout the remainder of the paper, let $n,m$ be positive integers and $r$ a nonnegative integer with $r\leq\min\{n,m\}$. As in the introduction, ${\mathcal{Z}}_{n,m,r}$ is the finite locus of rook placements of size $r$ on the $n\times m$ board. We can view ${\mathcal{Z}}_{n,m,r}\subseteq{\mathrm{Mat}}_{n\times m}({\mathbb{C}})$ as the locus of $n\times m$ complex matrices defined by the following conditions:
  • Every entry in the matrix is 0 or 11
  • There is at most one nonzero entry in each row or column
  • There are $r$ ones in total.
Our object of study is the upper rook placements, defined as
(3.1) ${\mathcal{UZ}}_{n,m,r}\coloneqq\bigsqcup_{r^{\prime}=r}^{\min\{m,n\}}{\mathcal{Z}}_{n,m,r^{\prime}}\subseteq{\mathrm{Mat}}_{n\times m}({\mathbb{C}})$
the union of all rook placements of size at least $r$. To study the orbit harmonics quotient $R({\mathcal{Z}}_{n,m,r})$, we define the following ideal, which we later prove to be ${\mathrm{gr}}\,{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$.
Definition 3.1.
Let ${\mathbf{x}}_{n\times m}$ be an $n\times m$ matrix of variables $(x_{i,j})_{1\leq i\leq n,1\leq j\leq m}$ and consider the polynomial ring ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]$ over these variables. Let $I_{n,m,r}\subseteq{\mathbb{C}}[{\mathbf{x}}_{n\times m}]$ be the ideal generated by
  • any product $x_{i,j}\cdot x_{i,j^{\prime}}\,(1\leq i\leq n;\,1\leq j,j^{\prime}\leq m)$ of variables in the same row,
  • any product $x_{i,j}\cdot x_{i^{\prime},j}\,(1\leq i,i^{\prime}\leq n;\,1\leq j\leq m)$ of variables in the same column,
  • any product $\prod_{k=1}^{n-r+1}\big(\sum_{j=1}^{m}x_{i_{k},j}\big)$ of nr+1n-r+1 distinct row sums for $1\leq i_{1}<\cdots<i_{n-r+1}\leq n$, and
  • any product $\prod_{k=1}^{m-r+1}\big(\sum_{i=1}^{n}x_{i,j_{k}}\big)$ of mr+1m-r+1 distinct column sums for $1\leq j_{1}<\cdots<j_{m-r+1}\leq m$.
To study this ideal, we will also need the following definition, which associates a monomial with any rook placement:
Definition 3.2.
For any rook placement $\mathcal{R}$} on the $n\times m$ board, the monomial ${\mathfrak{m}}({\mathcal{R}})\in{\mathbb{C}}[{\mathbf{x}}_{n\times m}]$ is defined to be
${\mathfrak{m}}({\mathcal{R}})\coloneqq\prod_{(i,j)\in{\mathcal{R}}}x_{i,j}.$
In the following subsection, we first handle the special case $r=0$, which prepares the reader for the more technical Section 3.2.
3.1. $r=0$ case
For convenience, write ${\mathcal{Z}}_{n,m}={\mathcal{UZ}}_{n,m,0}$ for the set of all rook placements on the $n\times m$ board. Let $I_{n,m}=I_{n,m,0}$ be the ideal generated by
  • any product $x_{i,j}\cdot x_{i,j^{\prime}}$ for $1\leq i\leq n$ and $1\leq j,j^{\prime}\leq m$ of variables in the same row, and
  • any product $x_{i,j}\cdot x_{i^{\prime},j}$ for $1\leq i,i^{\prime}\leq n$ and $1\leq j\leq m$ of variables in the same column.
We start by giving a spanning set of the vector space ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}$:
Lemma 3.3.
The family of monomials $\{{\mathfrak{m}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{Z}}_{n,m}\}$ descends to a spanning set of the vector space ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}$.
Proof.
Note that the generators $x_{i,j}\cdot x_{i,j^{\prime}}$ and $x_{i,j}\cdot x_{i^{\prime},j}$ of $I_{n,m}$ above imply that all the monomials containing variables from the same column or row of ${\mathbf{x}}_{n\times m}$ vanish in the quotient ring ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}$. Consequently, the monomials with no two variables from the same row or column span ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}$. ∎
We next strengthen Lemma 3.3 to yield a basis of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}$.
Proposition 3.4.
The family of monomials $\{{\mathfrak{m}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{Z}}_{n,m}\}$ descends to a $\mathbb{C}$}-linear basis of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}$. Furthermore, we have ${\mathrm{gr}}{\mathbf{I}}({\mathcal{Z}}_{n,m})=I_{n,m}$.
Proof.
Since a rook placement contains no rooks in the same row or column, it is clear that $x_{i,j}\cdot x_{i,j^{\prime}}\in{\mathbf{I}}({\mathcal{Z}}_{n,m})$ and $x_{i,j}\cdot x_{i^{\prime},j}\in{\mathbf{I}}({\mathcal{Z}}_{n,m})$. We thus have $x_{i,j}\cdot x_{i,j^{\prime}}\in{\mathrm{gr}}{\mathbf{I}}({\mathcal{Z}}_{n,m})$ and $x_{i,j}\cdot x_{i^{\prime},j}\in{\mathrm{gr}}{\mathbf{I}}({\mathcal{Z}}_{n,m})$. This implies that $I_{n,m}\subseteq{\mathrm{gr}}{\mathbf{I}}({\mathcal{Z}}_{n,m})$, which naturally induces a linear surjection
${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}\twoheadrightarrow R({\mathcal{Z}}_{n,m}).$
We now compare dimensions. Lemma 3.3 indicates that
$\dim_{\mathbb{C}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m})\leq|{\mathcal{Z}}_{n,m}|=\dim_{\mathbb{C}}(R({\mathcal{Z}}_{n,m})),$
while the surjection above implies $\dim_{\mathbb{C}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m})\geq\dim_{\mathbb{C}}(R({\mathcal{Z}}_{n,m}))$. Hence, all inequalities are equalities, and the surjection above is a linear isomorphism. This forces $I_{n,m}={\mathrm{gr}}{\mathbf{I}}({\mathcal{Z}}_{n,m})$ and ensures that the spanning set $\{{\mathfrak{m}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{Z}}_{n,m}\}$ in Lemma 3.3 is a basis. ∎
The monomial basis from Proposition 3.4 inherits the ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-action on ${\mathcal{Z}}_{n,m}$, allowing us to compute the graded module structure of $R({\mathcal{Z}}_{n,m})={\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}$.
Proposition 3.5.
The graded Frobenius image of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}$ is given by
${\mathrm{grFrob}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m};q)=\sum_{d=0}^{\min\{m,n\}}q^{d}\cdot\bigg(\sum_{\mu\vdash d}(s_{\mu}\cdot h_{n-d})\otimes(s_{\mu}\cdot h_{m-d})\bigg).$
Proof.
Note that the monomial basis in Proposition 3.4 is compatible with the ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-action on ${\mathcal{Z}}_{n,m}$. Therefore, we have an isomorphism of ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-modules:
$({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m})_{d}\cong{\mathbb{C}}[{\mathcal{Z}}_{n,m,d}]$
for $0\leq d\leq\min\{m,n\}$, and
$({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m})_{d}=\{0\}$
for $d>\min\{m,n\}$.
It suffices to study the module structure of ${\mathbb{C}}[{\mathcal{Z}}_{n,m,d}]$ for $0\leq d\leq\min\{m,n\}$. Consider ${\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d}$ (resp. ${\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d}$) as a subgroup of ${\mathfrak{S}}_{n}$ (resp. ${\mathfrak{S}}_{m}$) by letting ${\mathfrak{S}}_{d}$ and ${\mathfrak{S}}_{n-d}$ (resp. ${\mathfrak{S}}_{m-d}$) separately act on $\{1,\dots,d\}$ and $\{d+1,\dots,n\}$ (resp. $\{d+1,\cdots,m\}$). Thus, we embed $({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d})\times({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d})$ into ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$ as a subgroup. Note that $({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d})\times({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d})$ acts on ${\mathbb{C}}[{\mathfrak{S}}_{d}]$ by
$((g_{1},h_{1}),(g_{2},h_{2}))\cdot g=g_{1}gg_{2}^{-1}$
which restricts to the ${\mathfrak{S}}_{d}\times{\mathfrak{S}}_{d}$-action. Artin-Wedderburn Theorem gives this ${\mathfrak{S}}_{d}\times{\mathfrak{S}}_{d}$-module structure by
(3.2) ${\mathbb{C}}[{\mathfrak{S}}_{d}]\cong\bigoplus_{\mu\vdash d}V^{\mu}\otimes V^{\mu}$
where ${\mathfrak{S}}_{d}\times{\mathfrak{S}}_{d}$ acts on each $V^{\mu}\otimes V^{\mu}$ by $(g_{1},g_{2})\cdot(v_{1}\otimes v_{2})=(g_{1}\cdot v_{1})\otimes(g_{2}\cdot v_{2})$. Then we have that, as ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-modules,
${\mathbb{C}}[{\mathcal{Z}}_{n,m,d}]$ $\cong\mathrm{Ind}_{({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d})\times({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d})}^{{\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}}{\mathbb{C}}[{\mathfrak{S}}_{d}]$
$\cong\mathrm{Ind}_{({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d})\times({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d})}^{{\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}}\bigg(\bigoplus_{\mu\vdash d}(V^{\mu}\otimes V^{n-d})\otimes(V^{\mu}\otimes V^{m-d})\bigg)$
$\cong\bigoplus_{\mu\vdash d}\mathrm{Ind}_{({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d})\times({\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d})}^{{\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}}\big((V^{\mu}\otimes V^{n-d})\otimes(V^{\mu}\otimes V^{m-d})\big)$
$\cong\bigoplus_{\mu\vdash d}\Big(\mathrm{Ind}_{{\mathfrak{S}}_{d}\times{\mathfrak{S}}_{n-d}}^{{\mathfrak{S}}_{n}}(V^{\mu}\otimes V^{n-d})\Big)\otimes\Big(\mathrm{Ind}_{{\mathfrak{S}}_{d}\times{\mathfrak{S}}_{m-d}}^{{\mathfrak{S}}_{m}}(V^{\mu}\otimes V^{m-d})\Big)$
where the second isomorphism arises from Equation (3.2). Therefore, the Frobenius image of ${\mathbb{C}}[{\mathcal{Z}}_{n,m,d}]$ as an ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-module is given by
(3.3) ${\mathrm{Frob}}({\mathbb{C}}[{\mathcal{Z}}_{n,m,d}])=\sum_{\mu\vdash d}(s_{\mu}\cdot h_{n-d})\otimes(s_{\mu}\cdot h_{m-d})$
and hence
${\mathrm{Frob}}\big(({\mathbb{C}}[{\mathbf{x}}_{n\times m}])_{d}\big)=\sum_{\mu\vdash d}(s_{\mu}\cdot h_{n-d})\otimes(s_{\mu}\cdot h_{m-d}),$
completing our proof. ∎
3.2. Structure of $I_{n,m,r}$
In order to analyze the standard monomial basis of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ in Section 4 and the graded module structure of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ in Section 5, we start by studying strategic elements of the ideal $I_{n,m,r}$. On a first reading, it may be better to skip all the technical proofs in this section and accept Corollary 3.9 and Corollary 3.10 on faith.
For induction convenience, we introduce a new ideal $J_{n,m,r}\subseteq I_{n,m,r}$ by removing some generators of $I_{n,m,r}$ in Definition 3.1:
Definition 3.6.
Let $J_{n,m,r}\subseteq{\mathbb{C}}[{\mathbf{x}}_{n\times m}]$ be the ideal generated by
  • all products $x_{i,j}\cdot x_{i,j^{\prime}}$ for $1\leq i\leq n$ and $1\leq j,j^{\prime}\leq m$ of variables in the same row, and
  • all products $\prod_{k=1}^{m-r+1}\big(\sum_{i=1}^{n}x_{i,j_{k}}\big)$ of mr+1m-r+1 distinct column sums for $1\leq j_{1}<\cdots<j_{m-r+1}\leq m$.
We start by modifying the products of column sums in Definition 3.6 so they are easier to use:
Lemma 3.7.
If $m-r<t\leq m$ and $b_{1},\cdots,b_{t}\in[m]$ are distinct integers, then
$\sum_{\begin{subarray}{c}i_{1},\cdots,i_{t}\in[n]\\ \text{distinct}\end{subarray}}\prod_{j=1}^{t}x_{i_{j},b_{j}}\in J_{n,m,r}$
summing over all ordered sequences of distinct integers in $[n]$ of length $t$.
Proof.
Since any product of the form $x_{i,j}\cdot x_{i,j^{\prime}}$ belongs to $J_{n,m,r}$, we have that
$\sum_{\begin{subarray}{c}i_{1},\cdots,i_{t}\in[n]\\ \text{distinct}\end{subarray}}\prod_{j=1}^{t}x_{i_{j},b_{j}}\equiv\sum_{i_{1},\cdots,i_{t}\in[n]}\prod_{j=1}^{t}x_{i_{j},b_{j}}=\prod_{j=1}^{t}\sum_{i=1}^{n}x_{i,b_{j}}\mod{J_{n,m,r}}.$
Note that t>mrt>m-r. Then the generators of $J_{n,m,r}$ of the form $\prod_{k=1}^{m-r+1}\big(\sum_{i=1}^{n}x_{i,j_{k}}\big)$ imply that $\prod_{j=1}^{t}\sum_{i=1}^{n}x_{i,b_{j}}\in J_{n,m,r}$, completing the proof. ∎
We will then prove the most technical result in this section. As we have the $\mathbb{C}$}-algebra identification ${\mathbb{C}}[{\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}]\cong{\mathbb{C}}[{\mathfrak{S}}_{n}]\otimes_{\mathbb{C}}{\mathbb{C}}[{\mathfrak{S}}_{m}]$, we write $\eta_{p}\otimes 1$ for $\sum_{w\in{\mathfrak{S}}_{p}}(w,1)$.
Lemma 3.8.
Let $p$ and $d$ be two integers such that $0\leq d\leq\min\{m,n\}$ and $n+m-d-r<p\leq n$. Then, for any rook placement ${\mathcal{R}}\in{\mathcal{Z}}_{n,m,d}$, we have
$(\eta_{p}\otimes 1)\cdot{\mathfrak{m}}({\mathcal{R}})\in J_{n,m,r}.$
Proof.
We prove the lemma by inducting on $n$. The base case $n=1$ is trivial. Assume that Lemma 3.8 holds for any positive integer $n^{\prime}<n$, we prove it for $n$.
Write ${\mathfrak{m}}({\mathcal{R}})=\prod_{k=1}^{d}x_{a_{k},b_{k}}$ where $b_{1},\cdots,b_{d}\in[m]$ are distinct and $1\leq a_{1}<\cdots<a_{d}\leq n$. Since $p+d>n+m-r\geq n$, it follows that $p]\cap\{a_{1},\cdots,a_{d}\}\neq\varnothing$. Therefore, we can choose the largest $t\in[d]$ such that $a_{t}\in[p]$.
Claim: t>mrt>m-r.
In fact, the maximality of $t$ implies that $p]\cap\{a_{t+1},\cdots,a_{d}\}=\varnothing$, so we have that $p]\sqcup\{a_{t+1},\cdots,a_{d}\}\subseteq[n]$ and hence $p+d-t\leq n$. Consequently, $t\geq p+d-n>m-r$, which completes the proof of the claim.
Back to the proof of Lemma 3.8. We have that
$(\eta_{p}\otimes 1)\cdot{\mathfrak{m}}({\mathcal{R}})=(\eta_{p}\otimes 1)\cdot\prod_{k=1}^{d}x_{a_{k},b_{k}}=(p-t)!\cdot\Bigg(\prod_{\begin{subarray}{c}i_{1},\cdots,i_{t}\in[p]\\ \text{distinct}\end{subarray}}x_{i_{1},b_{1}}\cdots x_{i_{t},b_{t}}\Bigg)\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}$
$\equiv$ $(p-t)!\cdot\Bigg(-\sum_{\begin{subarray}{c}\text{distinct $i_{1},\cdots,i_{t}\in[n]$,}\\ \{i_{1},\cdots,i_{t}\}\not\subseteq[p]\end{subarray}}x_{i_{1},b_{1}}\cdots x_{i_{t},b_{t}}\Bigg)\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}$
$\equiv$ $(p-t)!\cdot\Bigg(-\sum_{\begin{subarray}{c}\text{distinct $i_{1},\cdots,i_{t}\in[n]\setminus\{a_{t+1},\cdots,a_{d}\}$,}\\ \{i_{1},\cdots,i_{t}\}\not\subseteq[p]\end{subarray}}x_{i_{1},b_{1}}\cdots x_{i_{t},b_{t}}\Bigg)\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}\mod{J_{n,m,r}}$
where the first congruence follows from the claim above, and the second from the relation $x_{i,j}\cdot x_{i,j^{\prime}}\in J_{n,m,r}$. Now it remains to show that
(3.4) $\Bigg(\sum_{\begin{subarray}{c}\text{distinct $i_{1},\cdots,i_{t}\in[n]\setminus\{a_{t+1},\cdots,a_{d}\}$,}\\ \{i_{1},\cdots,i_{t}\}\not\subseteq[p]\end{subarray}}x_{i_{1},b_{1}}\cdots x_{i_{t},b_{t}}\Bigg)\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}\in J_{n,m,r}.$
If the summation in (3.4) is empty, there is nothing to prove. Otherwise, we decompose the summation into sub-summations with smaller index sets as follows. Note that the summation in (3.4) requires that $\{i_{1},\cdots,i_{t}\}\not\subseteq[p]$, so we assume, without loss of generality, that $\{i_{1},\cdots,i_{t}\}\setminus[p]=\{i_{t-s+1},\cdots,i_{t}\}$ where $0<s<t$. (Here, we have $s<t$ because $\{i_{1},\cdots,i_{t}\}\cap[p]\neq\varnothing$. Otherwise, $\{i_{1},\cdots,i_{t}\}\sqcup[p]\sqcup\{a_{t+1},\cdots,a_{d}\}\subseteq[n]$ indicates that $t+p+d-t\leq n$, i.e. $p\leq n-d$. However, $p>n+m-d-r\geq n-d$, a contradiction.) Summing over such $i_{1},\cdots,i_{t}$ for fixed $i_{t-s+1}=c_{t-s+1},\cdots,i_{t}=c_{t}$, we obtain a sub-summation of the summation in (3.4). It suffices to show (3.4) after replacing its summation with these sub-summations, since adding them up yields (3.4) itself. Consequently, it remains to show that
$\Bigg(\sum_{\begin{subarray}{c}i_{1},\cdots,i_{t-s}\in[p]\setminus\{a_{t+1},\cdots,a_{d}\}\\ \text{distinct}\end{subarray}}x_{i_{1},b_{1}}\cdots x_{i_{t-s},b_{t-s}}\Bigg)\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}\in J_{n,m,r}.$
Recall that the choice of $t$ implies that $p]\cap\{a_{t+1},\cdots,a_{d}\}=\varnothing$, so it remains to show that
(3.5) $\Bigg(\sum_{\begin{subarray}{c}i_{1},\cdots,i_{t-s}\in[p]\\ \text{distinct}\end{subarray}}x_{i_{1},b_{1}}\cdots x_{i_{t-s},b_{t-s}}\Bigg)\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}\in J_{n,m,r}.$
Note that this expression possesses the following form:
(3.6) $c\cdot((\eta_{p}\otimes 1)\cdot{\mathfrak{m}}({\mathcal{R}}^{\prime}))\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}$
where $c\in{\mathbb{C}}$, and $\mathcal{R}$$^{\prime}$ is a rook placement of size $d^{\prime}=d-s$ on the sub-board $([n]\setminus\{c_{t-s+1},\cdots,c_{t}\})\times([m]\setminus\{b_{t-s+1},\cdots,b_{t}\})$ of the original board $n]\times[m]$. For the sake of induction, consider restricting Definition 3.6 to a smaller variable set:
  • $n^{\prime}=n-s$, $m^{\prime}=m-s$, and $r^{\prime}=r-s$;
  • the restricted variable matrix given by
    ${\mathbf{x}}_{n^{\prime},m^{\prime}}^{\prime}\coloneqq(x_{i,j})_{i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\},j\in[m]\setminus\{b_{t-s+1},\cdots,b_{t}\}};$
  • the restricted counterpart $J^{\prime}_{n^{\prime},m^{\prime},r^{\prime}}\subseteq{\mathbb{C}}[{\mathbf{x}}]$ of $J_{n,m,r}$ generated by
    • any product $x_{i,j}\cdot x_{i,j^{\prime}}$ for $i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\}$ and $j,j^{\prime}\in[m]\setminus\{b_{t-s+1},\cdots,x_{t}\}$, and
    • any product $\prod_{k=1}^{m^{\prime}-r^{\prime}+1}(\sum_{i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\}}x_{i,j_{k}})$ for distinct $j_{1},\cdots,j_{m^{\prime}-r^{\prime}+1}\in[m]\setminus\{b_{t-s+1},\cdots,b_{t}\}$.
Note that $p>n+m-d-r=(n-s)+(m-s)-(d-s)-(r-s)=n^{\prime}+m^{\prime}-d^{\prime}-r^{\prime}$, and hence the induction assumption reveals that $(\eta_{p}\otimes 1)\cdot{\mathfrak{m}}({\mathcal{R}}^{\prime})\in J_{n^{\prime},m^{\prime},r^{\prime}}^{\prime}$. Therefore, in order to show (3.5) using Expression (3.6), it suffices to show that
$J_{n^{\prime},m^{\prime},r^{\prime}}^{\prime}\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}\subseteq J_{n,m,r}.$
We verify this containment using the generators of $J_{n^{\prime},m^{\prime},r^{\prime}}^{\prime}$ above. For any product $x_{i,j}\cdot x_{i,j^{\prime}}$ for $i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\}$ and $j,j^{\prime}\in[m]\setminus\{b_{t-s+1},\cdots,x_{t}\}$, it is clear that $x_{i,j}\cdot x_{i,j^{\prime}}\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}\in J_{n,m,r}$. For any product $\prod_{k=1}^{m^{\prime}-r^{\prime}+1}(\sum_{i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\}}x_{i,j_{k}})$ for distinct $j_{1},\cdots,j_{m^{\prime}-r^{\prime}+1}\in[m]\setminus\{b_{t-s+1},\cdots,b_{t}\}$, $m^{\prime}-r^{\prime}+1=(m-s)-(r-s)+1=m-r+1$ indicates that
$\prod_{k=1}^{m^{\prime}-r^{\prime}+1}\Bigg(\sum_{i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\}}x_{i,j_{k}}\Bigg)\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}$
$=$ $\prod_{k=1}^{m-r+1}\Bigg(\sum_{i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\}}x_{i,j_{k}}\Bigg)\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}$
$\equiv$ $\prod_{k=1}^{m-r+1}\Bigg(\sum_{i\in[n]}x_{i,j_{k}}\Bigg)\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}\equiv 0\mod{J_{n,m,r}}$
which completes the proof. ∎
Corollary 3.9.
Let $0\leq d\leq\min\{m,n\}$ be an integer. For any integer $n+m-d-r<p\leq n$ and any rook placement ${\mathcal{R}}\in{\mathcal{Z}}_{n,m,d}$, we have that
$(\eta_{p}\otimes 1)\cdot{\mathfrak{m}}({\mathcal{R}})\in I_{n,m,r}.$
Similarly, for any integer $n+m-d-r<\tilde{p}\leq m$, we have that
$(1\otimes\eta_{\tilde{p}})\cdot{\mathfrak{m}}({\mathcal{R}})\in I_{n,m,r}.$
Proof.
Since $J_{n,m,r}\subseteq I_{n,m,r}$, Lemma 3.8 immediately indicates that $(\eta_{p}\otimes 1)\cdot{\mathfrak{m}}({\mathcal{R}})\in I_{n,m,r}$. As we did for Lemma 3.8, we can similarly define an ideal $K_{n,m,r}\subseteq I_{n,m,r}$ by interchanging the roles of rows and columns in Definition 3.6. This construction gives $(1\otimes\eta_{\tilde{p}})\cdot{\mathfrak{m}}({\mathcal{R}})\in K_{n,m,r}$, which implies that $(1\otimes\eta_{\tilde{p}})\cdot{\mathfrak{m}}({\mathcal{R}})\in I_{n,m,r}$. ∎
Corollary 3.10.
Let $0\leq d\leq\min\{m,n\}$ be an integer. For any set $S\subseteq[n]$ and $T\subseteq[m]$ such that |S|,|T|>n+mdr|S|,|T|>n+m-d-r, let ${\mathfrak{S}}_{S}\subseteq{\mathfrak{S}}_{n}$ (resp. ${\mathfrak{S}}_{T}\subseteq{\mathfrak{S}}_{m}$) be the symmetric group of $S$ (resp. $T$). Then, for any rook placement ${\mathcal{R}}\in{\mathcal{Z}}_{n,m,d}$, we have
$\sum_{w\in{\mathfrak{S}}_{S}}(w,1)\cdot{\mathfrak{m}}({\mathcal{R}})\in I_{n,m,r}$
and
$\sum_{w\in{\mathfrak{S}}_{T}}(1,w)\cdot{\mathfrak{m}}({\mathcal{R}})\in I_{n,m,r}.$
Proof.
It follows from Corollary 3.9 immediately, because $I_{n,m,r}$ is closed under the action of ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$. ∎
4. Hilbert series and standard monomial basis
4.1. Viennot Shadow
We begin this section by describing Viennot’s shadow line construction [16]. Let $w=[w(1),w(2),\dots,w(n)]\in{\mathfrak{S}}_{n}$ be a permutation, and we draw its diagram on an $n\times n$ grid by putting a point at $(i,w(i))$ for $i\in[n]$. For example, the permutation $w=[6,3,5,7,1,2,8,4]\in{\mathfrak{S}}_{8}$ has the digram shown below.
Place an imaginary light source at the bottom left corner of the diagram, shining northeast. Each point $(i,w(i))$ in the diagram of $w$ blocks the regions to its north and to its east. We take the boundary of the shaded region and call it the first shadow line $L_{1}$ (see the left diagram below). Removing all points on the first shadow line and iterating this process yields the second shadow line $L_{2}$, third shadow line $L_{3}$, and so on (see the center diagram below).
$L_{1}$$L_{1}$$L_{2}$$L_{3}$$L_{4}$$L_{1}$$L_{2}$$L_{3}$$L_{4}$
[…]
Definition 4.1.
Given a permutation $w\in{\mathfrak{S}}_{n}$, the shadow set ${\mathcal{S}}(w)$ of $w$ is the collection of points that lie on the northeast corners of the shadow lines.
[…]
This iterated process produces the iterated shadow set ${\mathcal{S}}({\mathcal{S}}(w))$, which consists of a single point as drawn in red above. Repeating this process, we obtain a single shadow line and an empty shadow set,
from which we conclude that $P(w)$ and $Q(w)$ are respectively given by
11 88 33 77 66
and   11 77 22 88 55
Remark 4.2.
From the construction, it is immediate that the shadow set ${\mathcal{S}}(w)$ of any permutation $w$ is a rook placement. However, not all rook placements are shadow sets of permutations. In  [13], Rhoades described an algorithm to determine whether a rook placement is the shadow set of some permutation. In the next section, we present a generalized version of this algorithm.
4.2. Extended shadow sets
Recall that ${\mathcal{UZ}}_{n,m,r}$ is the set of rook placements with at least $r$ rooks. Given a rook placement ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}$, we define an algorithm that extends $\mathcal{R}$} to a permutation ${\mathcal{EX}}({\mathcal{R}})\in{\mathfrak{S}}_{n+m-r}$. The shadow set of this permutation, denoted by ${\mathcal{ES}}(R)$, plays a key role in analyzing the orbit harmonics quotient $R({\mathcal{UZ}}_{n,m,r})$.
To start, let ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}$ with $r^{\prime}\geq r$ rooks. Draw the diagram of the rook placement on the $n]\times[m]$ grid, with a point at $(i,j)$ if $(i,j)\in R$. Below is the diagram of ${\mathcal{R}}=\{(2,3),(3,4),(5,2),(8,5)\}\in{\mathcal{UZ}}_{8,6,2}$.
1122334455667788112233445566
We extend the diagram to an $(n+m-r)\times(n+m-r)$ grid by adding (mr)(m-r) columns on the left and (nr)(n-r) rows on the bottom. The newly added rows and columns are drawn in red in the diagram below. Label the added rows as $\{\overline{1},\overline{2},\dots,\overline{n-r}\}$ from bottom to top and label the added columns as $\{\overline{1},\overline{2},\dots,\overline{m-r}\}$ from left to right, as shown in the diagram.
$\overline{1}$$\overline{2}$$\overline{3}$$\overline{4}$$\overline{5}$$\overline{6}$$\overline{1}$$\overline{2}$$\overline{3}$$\overline{4}$1122334455667788112233445566
For $i\leq r^{\prime}-r$, we put a point at $(\overline{i},\overline{i})$, as drawn in black in the diagram below. Finally, let $i_{1}<i_{2}<\dots<i_{n-r^{\prime}}$ be the indices of columns that do not contain a rook, and let $j_{1}<j_{2}<\dots<j_{m-r^{\prime}}$ be the indices of rows that do not contain a rook. For $1\leq l\leq m-r^{\prime}$, add a point at $(\overline{r^{\prime}-r+l},j_{l})$; for $1\leq k\leq n-r^{\prime}$, add a point at $(i_{k},\overline{r^{\prime}-r+k})$. These added points are drawn in red in the diagram below.
$\overline{1}$$\overline{2}$$\overline{3}$$\overline{4}$$\overline{5}$$\overline{6}$$\overline{1}$$\overline{2}$$\overline{3}$$\overline{4}$1122334455667788112233445566
It can be easily checked that the diagram obtained this way has exactly one point in each row and each column, hence forming a permutation in ${\mathfrak{S}}_{n+m-r}$. We write ${\mathcal{EX}}({\mathcal{R}})$ for this extended permutation arising from $\mathcal{R}$}, identified with the diagram on the extended board.
Remark 4.3.
There is another way to define the extended permutation ${\mathcal{EX}}({\mathcal{R}})$. Define an ordering of row and column indices by
  • if $i<j$ then $\overline{i}<\overline{j}$
  • $\overline{i}<j$ for all $i,j$.
${\mathcal{EX}}({\mathcal{R}})$ is the unique extension of $\mathcal{R}$} to a permutation in the $(n+m-r)\times(n+m-r)$ board such that the vertical coordinates of the points in the first mrm-r columns form an increasing pattern and the horizontal coordinates of the points in the first nrn-r rows form an increasing pattern.
Apply Viennot’s shadow line construction to ${\mathcal{EX}}(R)$. From the construction (and by Remark 4.3), we have that the shadow lines $L_{1},L_{2},\dots,L_{n-r}$ will have infinite horizontal coordinates $\overline{1},\overline{2},\dots,\overline{n-r}$, and the shadow lines $L_{1},L_{2},\dots,L_{m-r}$ will have infinite vertical coordinates $\overline{1},\overline{2},\dots,\overline{m-r}$. Let ${\mathcal{ES}}({\mathcal{R}})$ be the shadow set of ${\mathcal{EX}}({\mathcal{R}})$, we call it the extended shadow set of $\mathcal{R}$}. See the diagram for an example.
[…]
The following lemma is an important fact about ${\mathcal{ES}}({\mathcal{R}})$.
Lemma 4.4.
Let ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}$, we have that ${\mathcal{ES}}({\mathcal{R}})$ is also a rook placement in $n]\times[m]$. That is, the shadow set of ${\mathcal{EX}}({\mathcal{R}})$ is contained in the $n\times m$ box in the upper right corner of the diagram, with column indices in $[n]$ and row indices in $[m]$.
Proof.
Let $(P,Q)$ be the pair of tableaux corresponding to ${\mathcal{EX}}({\mathcal{R}})$ in the Schensted correspondence. We adopt the notion in Remark 4.3 so the content of $P$ is $\{\overline{1},\overline{2},\dots,\overline{n-r},1,2,\dots m\}$ and the content of $Q$ is $\{\overline{1},\overline{2},\dots,\overline{m-1},1,2,\dots n\}$. According to Viennot’s construction in  [16], we have that $\overline{1},\overline{2},\dots,\overline{n-r}$ are in the first row of $P$ as they are the coordinates of the infinite horizontal rays of shadow lines $L_{1},L_{2},\dots,L_{n-r}$. Similarly, $\overline{1},\overline{2},\dots,\overline{m-r}$ are in the first row of $Q$. As the horizontal coordinates of points in the shadow set are entries in $P$ beyond the first row, and the vertical coordinates of points in the shadow set are entries in $Q$ beyond the shadow set, we conclude that ${\mathcal{ES}}({\mathcal{R}})={\mathcal{S}}({\mathcal{EX}}({\mathcal{R}}))\subseteq[n]\times[m]$. ∎
Given a rook placement ${\mathcal{R}}\subseteq[n]\times[m]$, the following proposition characterizes whether $\mathcal{R}$} is the extended shadow set of some other rook placement $\tilde{{\mathcal{R}}}$ with size at least $r$.
Proposition 4.5.
Let $\mathcal{R}$} be a rook placement on the $n\times m$ board and apply the Viennot shadow construction to $\mathcal{R}$}, yielding shadow lines $L_{1},\cdots,L_{p}$. Define two sequences $(x_{i})_{i=1}^{n}$ and $(y_{j})_{j=1}^{m}$ over the alphabet $\{-1,0,1\}$ by
[…]
and
[…]
Proof.
[…]
4.3. Extended shadow monomials and spanning
Our next goal is to obtain as small a spanning set of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ as possible using the combinatorics of the previous section.
Lemma 4.6.
The family of monomials $\{{\mathfrak{m}}({\mathcal{R}})\,:\,{\mathcal{R}}\in\bigsqcup_{d=0}^{\min\{m,n\}}{\mathcal{Z}}_{n,m,d}\}$ descends to a spanning set of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$.
Proof.
This follows immediately since $I_{n,m,r}$ contains all the products $x_{i,j}\cdot x_{i,j^{\prime}}$ and $x_{i,j}\cdot x_{i^{\prime},j}$ of variables in the same row or column. ∎
However, the spanning set given by Lemma 4.6 is far from a basis. To extract a basis from this spanning set, we introduce a class of special monomial orders, namely the diagonal monomial orders. Recall that the lexicographical order in an ordered set of variables $y_{1}>y_{2}>\cdots>y_{N}$ is given by $y_{1}^{a_{1}}\cdots y_{N}^{a_{N}}<y_{1}^{b_{1}}\cdots y_{N}^{b_{N}}$ if and only if there exists an integer $1\leq M\leq N$ such that $a_{i}=b_{i}$ for $1\leq i<M$ and $a_{M}<b_{M}$.
Definition 4.7.
Consider a lexicographical order in an ordered set of variables ${\mathbf{x}}_{n\times m}$. If the order of the variable set satisfies $x_{i,j}>x_{i^{\prime},j^{\prime}}$ whenever $i+j<i^{\prime}+j^{\prime}$, then we call the lexicographical order a diagonal monomial order.
It is easy to construct a diagonal monomial order. For instance, the variable order
$x_{1,1}>x_{2,1}>x_{1,2}>x_{3,1}>x_{2,2}>x_{1,3}>\cdots>x_{n,m-1}>x_{n-1,m}>x_{n,m}$
induces a diagonal monomial order.
From now on, we fix a diagonal monomial order $_{\text{Diag}}$ on the set of monomials in ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]$.
Definition 4.8.
Let ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}$. The extended shadow monomial ${\mathfrak{es}}({\mathcal{R}})\in{\mathbb{C}}[{\mathbf{x}}_{n\times m}]$ is defined as ${\mathfrak{es}}({\mathcal{R}})\coloneqq{\mathfrak{m}}({\mathcal{ES}}({\mathcal{R}}))$.
We now prove the last technical lemma to obtain a smaller spanning set.
Lemma 4.9.
For $0\leq d\leq\min\{m,n\}$ and a rook placement ${\mathcal{R}}\in{\mathcal{Z}}_{n,m,d}$ that is not the extended shadow set of any rook placement in ${\mathcal{UZ}}_{n,m,r}$, we have
${\mathfrak{m}}({\mathcal{R}})\in\mathrm{span}\{{\mathfrak{m}}({\mathcal{R}}^{\prime})\,:\,{\mathcal{R}}^{\prime}\in{\mathcal{Z}}_{n,m,d},\,{\mathfrak{m}}({\mathcal{R}}^{\prime})<_{\text{Diag}}{\mathfrak{m}}({\mathcal{R}})\}+I_{n,m,r}.$
Proof.
Applying Viennot’s shadow construction to $\mathcal{R}$}, we obtain shadow lines $L_{1},\cdots,L_{p}$, together with sequences $(x_{i})_{i=1}^{n}$ and $(y_{j})_{j=1}^{m}$ as in Proposition 4.5. Because $\mathcal{R}$} is not the extended shadow set of any rook placements in ${\mathcal{UZ}}_{n,m,r}$, Proposition 4.5 implies that either $\sum_{i=1}^{k}x_{i}>m-r$ for some $1\leq k\leq n$ or $\sum_{j=1}^{l}y_{j}>n-r$ for some $1\leq l\leq m$.
Without loss of generality, we assume that $\sum_{i=1}^{k}x_{i}>m-r$ for some $1\leq k\leq n$ and further require $k$ to be the smallest integer with this property. Then $x_{k}>0$ and thus $x_{k}=1$, implying that some shadow line $L_{t}$ has a vertical ray at $x=k$. We define a size $t$ subset ${\mathcal{R}}_{1}=\{(i_{1},j_{1}),\cdots,(i_{t},j_{t})\}\subseteq{\mathcal{R}}$ as follows. Start at the vertical ray of $L_{t}$ and go south. Let $(i_{t},j_{t})$ be the first rook of $\mathcal{R}$} encountered (in particular, we have $i_{t}=k$). Then go west from $(i_{t},j_{t})$ until we encounter a vertical segment of the shadow line $L_{t-1}$ and turn left. Go south until we encounter a rook $(i_{t-1},j_{t-1})\in{\mathcal{R}}$. Repeating this process, we obtain a sequence of rooks $(i_{t},j_{t}),(i_{t-1},j_{t-1}),\cdots,(i_{1},j_{1})$ sequentially, such that $(i_{u},j_{u})$ is on the shadow line $L_{u}$ ($u=1,\cdots,t$), with $i_{1}<i_{2}<\cdots<i_{t}$, and $j_{1}<j_{2}<\cdots<j_{t}$. Let ${\mathcal{R}}_{2}={\mathcal{R}}\setminus{\mathcal{R}}_{1}$.
For example, let $n=m=8$ and $r=7$. Let ${\mathcal{R}}=\{(2,2),(3,5),(4,3),(5,6),(6,1),(7,7)\}\in{\mathcal{Z}}_{8,8,6}$ be a rook placement on the $8\times 8$ board. We mark all rooks of $\mathcal{R}$} using red points and draw black shadow lines $L_{1},\cdots,L_{4}$. The sequence $(x_{i})_{i=1}^{8}$ is 1,1,1,0,1,0,1,1-1,1,1,0,1,0,1,-1 where the smallest $k$ such that $\sum_{i=1}^{k}x_{i}>m-r$ is $k=5$. Since $x=5$ corresponds to the vertical ray of the shadow line $L_{3}$, we start from this vertical ray and go south, then west, and so on, yielding $\{(i_{1},j_{1}),(i_{2},j_{2}),(i_{3},j_{3})\}$. Our track is shown in blue as follows.
xxyy
Consequently we obtain $(i_{1},j_{1})=(2,2)$, $(i_{2},j_{2})=(3,5)$, and $(i_{3},j_{3})=(5,6)$, so
${\mathcal{R}}_{1}=\{(2,2),(3,5),(5,6)\}$
and
${\mathcal{R}}_{2}={\mathcal{R}}\setminus\{(2,2),(3,5),(5,6)\}=\{(4,3),(6,1),(7,7)\}$
in the example above.
Back to our proof. Define $S\subseteq[n]$ by
$S=\{i_{1},i_{2},\cdots,i_{t}\}\sqcup\{i\in\{i_{t}+1,\cdots,n\}\,:\,\text{the vertical line $x=i$ does not meet ${\mathcal{R}}$}\}.$
Claim: |S|>n+mdr|S|>n+m-d-r.
In fact, $\sum_{i=1}^{k}x_{i}>m-r$ indicates that
$t>|\{i\in[i_{t}]\,:\,\text{the vertical line $x=i$ does not meet ${\mathcal{R}}$}\}|+m-r$
and hence
$|S|=$ $t+|\{i\in\{i_{t}+1,\cdots,n\}\,:\,\text{the vertical line $x=i$ does not meet ${\mathcal{R}}$}\}|$
$>$ $|\{i\in[i_{t}]\,:\,\text{the vertical line $x=i$ does not meet ${\mathcal{R}}$}\}|+m-r+$
$|\{i\in\{i_{t}+1,\cdots,n\}\,:\,\text{the vertical line $x=i$ does not meet ${\mathcal{R}}$}\}|$
$=$ $|\{i\in[n]\,:\,\text{the vertical line $x=i$ does not meet ${\mathcal{R}}$}\}|+m-r$
$=$ $n-d+m-r=n+m-d-r$
which completes the proof of the claim above.
Then Corollary 3.10 reveals that
(4.1) $\sum_{w\in{\mathfrak{S}}_{S}}(w,1)\cdot{\mathfrak{m}}({\mathcal{R}})\in I_{n,m,r}.$
Note that ${\mathfrak{m}}({\mathcal{R}})={\mathfrak{m}}({\mathcal{R}}_{1})\cdot{\mathfrak{m}}(R_{2})$ where for all $w\in{\mathfrak{S}}_{S}$ the action of $(w,1)$ does not change ${\mathfrak{m}}({\mathcal{R}}_{2})$. Therefore, we have that
$\sum_{w\in{\mathfrak{S}}_{S}}(w,1)\cdot{\mathfrak{m}}({\mathcal{R}})=\bigg(\sum_{w\in{\mathfrak{S}}_{S}}(w,1)\cdot{\mathfrak{m}}({\mathcal{R}}_{1})\bigg)\cdot{\mathfrak{m}}({\mathcal{R}}_{2})$
where ${\mathfrak{m}}({\mathcal{R}}_{1})$ is the initial monomial of $\sum_{w\in{\mathfrak{S}}_{S}}(w,1)\cdot{\mathfrak{m}}({\mathcal{R}}_{1})$ under the monomial order $_{\text{Diag}}$, indicating that ${\mathfrak{m}}({\mathcal{R}}_{1})\cdot{\mathfrak{m}}({\mathcal{R}}_{2})={\mathfrak{m}}({\mathcal{R}})$ is the initial monomial of $\sum_{w\in{\mathfrak{S}}_{S}}(w,1)\cdot{\mathfrak{m}}({\mathcal{R}})$. As a result, the initial monomial of the expression on the left-hand side of Equation (4.1) is ${\mathfrak{m}}({\mathcal{R}})$, finishing the proof. ∎
Now we are ready to provide a smaller spanning set than Lemma 4.6. In the next section, we will see that this spanning set is actually the standard monomial basis of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ with respect to $_{\text{Diag}}$
Lemma 4.10.
The family of monomials $\{{\mathfrak{es}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}\}$ descends to a spanning set of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$.
Proof.
By Lemma 4.6, it suffices to show that for all the rook placements $\mathcal{R}$} on the $n\times m$ board we have that
${\mathfrak{m}}({\mathcal{R}})\in\mathrm{span}\{{\mathfrak{es}}({\mathcal{R}}^{\prime})\,:\,{\mathcal{R}}^{\prime}\in{\mathcal{UZ}}_{n,m,r}\}+I_{n,m,r}.$
[…]
Remark 4.11.
In fact, the proof of Lemma 4.10 gives a slightly stronger result than Lemma 4.10 itself: For each rook placement $\mathcal{R}$} on the $n\times m$ board, we have that
${\mathfrak{m}}({\mathcal{R}})\in\mathrm{span}\{{\mathfrak{es}}({\mathcal{R}}^{\prime})\,:\,{\mathcal{R}}^{\prime}\in{\mathcal{UZ}}_{n,m,r},\,\,{\mathfrak{es}}({\mathcal{R}}^{\prime})\leq_{\text{Diag}}{\mathfrak{m}}({\mathcal{R}})\}+I_{n,m,r}.$
4.4. Standard monomial basis and Hilbert series
As promised, we strengthen Lemma 4.10 to obtain the standard monomial basis of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$.
Theorem 4.12.
The family of monomials $\{{\mathfrak{es}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}\}$ descends to the standard monomial basis of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$. Furthermore, we have ${\mathrm{gr}}{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})=I_{n,m,r}$.
Remark 4.13.
Consider the case $r=0$. Recall that Proposition 3.4 provides a basis $\{{\mathfrak{m}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{Z}}_{n,m}\}$ which contains the standard monomial basis $\{{\mathfrak{es}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{UZ}}_{n,m,0}={\mathcal{Z}}_{n,m}\}$ given by Theorem 4.12. Therefore, we have
$\{{\mathfrak{m}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{Z}}_{n,m}\}=\{{\mathfrak{es}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{UZ}}_{n,m,0}={\mathcal{Z}}_{n,m}\},$
indicating that
${\mathcal{Z}}_{n,m}=\{{\mathcal{ES}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{Z}}_{n,m}\}.$
In fact, this fact is also immediate from Proposition 4.5.
Proof of Theorem 4.12.
Claim: $I_{n,m,r}\subseteq{\mathrm{gr}}{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$.
In fact, for any rook placement ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}$, each column (resp. row) of $n]\times[m]$ has at most one rook. Consequently, for all $1\leq i\leq n$ and $1\leq j,j^{\prime}\leq m$ we have $x_{i,j}\cdot x_{i,j^{\prime}}\in{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$ and thus $x_{i,j}\cdot x_{i,j^{\prime}}\in{\mathrm{gr}}{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$ (resp. for all $1\leq i,i^{\prime}\leq n$ and $1\leq j\leq m$ we have $x_{i,j}\cdot x_{i^{\prime},j}\in{\mathrm{gr}}{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$). Additionally, since $\mathcal{R}$} has at least $r$ rooks, any union of nr+1n-r+1 distinct columns of $n]\times[m]$ must meet $\mathcal{R}$}. Thus, for all $1\leq i_{1}<\cdots<i_{n-r+1}\leq n$ we have
$\prod_{k=1}^{n-r+1}\Bigg(\sum_{j=1}^{m}x_{i_{k},j}-1\Bigg)\in{\mathbf{I}}({\mathcal{UZ}}_{n,m,r}),$
which implies that
$\prod_{k=1}^{n-r+1}\Bigg(\sum_{j=1}^{m}x_{i_{k},j}\Bigg)\in{\mathrm{gr}}{\mathbf{I}}({\mathcal{UZ}}_{n,m,r}).$
Similarly, for all $1\leq j_{1}<\cdots<j_{m-r+1}\leq m$ we have
$\prod_{k=1}^{m-r+1}\Bigg(\sum_{i=1}^{n}x_{i,j_{k}}\Bigg)\in{\mathrm{gr}}{\mathbf{I}}({\mathcal{UZ}}_{n,m,r}).$
Then Definition 3.1 indicates the claim above.
The claim above yields a surjective $\mathbb{C}$}-linear map
${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}\twoheadrightarrow R({\mathcal{UZ}}_{n,m,r}),$
implying that $\dim_{\mathbb{C}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r})\geq\dim_{\mathbb{C}}(R({\mathcal{UZ}}_{n,m,r}))$. However, Lemma 4.10 indicates that $\dim_{\mathbb{C}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r})\leq|{\mathcal{UZ}}_{n,m,r}|=\dim_{\mathbb{C}}(R({\mathcal{UZ}}_{n,m,r}))$. Therefore, we conclude that
$\dim_{\mathbb{C}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r})=\dim_{\mathbb{C}}(R({\mathcal{UZ}}_{n,m,r})),$
so the spanning set $\{{\mathfrak{es}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}\}$ in Lemma 4.10 is indeed a basis of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$. Furthermore, the surjection ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}\twoheadrightarrow R({\mathcal{UZ}}_{n,m,r})$ above is actually a linear isomorphism, and hence
$I_{n,m,r}={\mathrm{gr}}{\mathbf{I}}({\mathcal{UZ}}_{n,m,r}).$
As we have shown that $\{{\mathfrak{es}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}\}$ descends to a basis of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$, it remains to show that $\{{\mathfrak{es}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}\}$ descends to the standard monomial basis of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$. Assume, for the sake of contradiction, that there exists some ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}$ such that ${\mathfrak{es}}({\mathcal{R}})$ is the initial monomial of some $f\in I_{n,m,r}$. Then we have that
(4.2) ${\mathfrak{es}}({\mathcal{R}})\equiv\sum_{\begin{subarray}{c}{\mathcal{R}}^{\prime}\in{\mathcal{Z}}_{n,m,|{\mathcal{R}}|}\\ {\mathfrak{m}}({\mathcal{R}}^{\prime})<{\mathfrak{es}}({\mathcal{R}})\end{subarray}}c_{{\mathcal{R}}^{\prime}}\cdot{\mathfrak{m}}({\mathcal{R}}^{\prime})\mod{I_{n,m,r}}$
where $c_{{\mathcal{R}}^{\prime}}\in{\mathbb{C}}$. However, Remark 4.11 indicates that the right-hand side of Equation 4.2 can be replaced with a linear combination of some extended shadow monomials
$\{{\mathfrak{es}}({\mathcal{R}}^{\prime})\,:\,{\mathcal{R}}^{\prime}\in{\mathcal{UZ}}_{n,m,r},\,\,{\mathfrak{es}}({\mathcal{R}}^{\prime})<_{\text{Diag}}{\mathfrak{es}}({\mathcal{R}})\},$
which means that extended shadow monomials $\{{\mathfrak{es}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}\}$ are linearly dependent in ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$. This yields a contradiction because we have shown the family of monomials $\{{\mathfrak{es}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}\}$ descends to a basis of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$. Therefore, the family of monomials $\{{\mathfrak{es}}({\mathcal{R}})\,:\,{\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}\}$ descends to the standard monomial basis of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$. ∎
Theorem 4.12 immediately gives the Hilbert series of $R({\mathcal{UZ}}_{n,m,r})={\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$. Given a permutation $w\in{\mathfrak{S}}_{n+m-r}$, a increasing subsequence of length $k$ is a sequence $1\leq i_{1}<\cdots<i_{k}\leq n+m-r$ of integers such that $w(i_{1})<\cdots<w(i_{k})$. We write
$\mathrm{lis}(w)\coloneqq\text{length of the longest increasing subsequence of $w$}.$
Corollary 4.14.
The Hilbert series of $R({\mathcal{UZ}}_{n,m,r})$ is
(4.3) ${\mathrm{Hilb}}(R({\mathcal{UZ}}_{n,m,r});q)=\sum_{{\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}}q^{n+m-r-\mathrm{lis}({\mathcal{EX}}({\mathcal{R}}))}.$
Proof.
By Theorem 4.12, it suffices to show that $\deg({\mathfrak{es}}({\mathcal{R}}))=n+m-r-\mathrm{lis}({\mathcal{EX}}({\mathcal{R}}))$ for ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}$, which is equivalent to
$\lvert{\mathcal{ES}}(w)\rvert=n+m-r-\mathrm{lis}({\mathcal{EX}}({\mathcal{R}})).$
Recall that ${\mathcal{ES}}({\mathcal{R}})={\mathcal{S}}({\mathcal{EX}}({\mathcal{R}}))$ is the shadow set of ${\mathcal{EX}}({\mathcal{R}})\in{\mathfrak{S}}_{n+m-r}$. Consequently, it remains to show that for all $w\in{\mathfrak{S}}_{n+m-r}$ we have
(4.4) $\lvert{\mathcal{S}}(w)\rvert=n+m-r-\mathrm{lis}(w).$
Let $k$ be the number of shadow lines of $w$ that we use to construct ${\mathcal{S}}(w)$. Write $w\mapsto(P,Q)$ according to the Schensted correspondence, where $P,Q$ have the same shape $\lambda$. The shadow line construction in Subsection 4.1 indicates that $k=\lambda_{1}$. In addition, Schensted correspondence possesses the standard property [15, Thm. 1] $\lambda_{1}=\mathrm{lis}(w)$, so we have $k=\mathrm{lis}(w)$ and thus Equation (4.4) holds. ∎
5. Module structure
We first provide some combinatorial results about symmetric functions. For convenience, we write $\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle F$ for the coefficient of $s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}$ in the Schur expansion of any $F\in\Lambda\otimes_{{\mathbb{C}}(q)}\Lambda$. Similarly, we write $\langle q^{i}\rangle f$ for the coefficient of $q^{i}$ in any generating function $f\in{\mathbb{C}}[[q]]$.
Lemma 5.1.
For $d,a,b,p,q\in{\mathbb{Z}}_{\geq 0}$, we consider
$F=\sum_{\mu\vdash d}\{s_{\mu}\cdot h_{a}\}_{\lambda_{1}=p}\otimes\{s_{\mu}\cdot h_{b}\}_{\lambda_{1}=q}\in\Lambda\otimes_{{\mathbb{C}}(q)}\Lambda$
and let $\lambda^{(1)}\vdash a+d$ and $\lambda^{(2)}\vdash b+d$ be partitions.
If all the following conditions hold:
  • $\lambda_{1}^{(1)}=p$,
  • $\lambda_{1}^{(2)}=q$, and
  • for all $i\geq 1$ we have $\min\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}\geq\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}$.
Then we have
$\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle F=\Big\langle q^{d-\sum_{i=1}^{\infty}\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}}\Big\rangle\Bigg(\prod_{i=1}^{\infty}\Bigg(\sum_{j=0}^{\min\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}-\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}}q^{j}\Bigg)\Bigg).$
Otherwise, we have
$\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle F=0.$
Proof.
If $s_{\lambda_{1}^{(1)}}\otimes s_{s_{\lambda_{1}^{(2)}}}$ appears in the Schur expansion of $F$, we must have $\lambda_{1}^{(1)}=p$ and $\lambda_{1}^{(2)}=q$ by the definition of $F$. Additionally, we must have $\min\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}\geq\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}$ for all $i\geq 1$, since both $\lambda^{(1)}$ and $\lambda^{(2)}$ arise from attaching a horizontal strip to the same partition $\mu$. In order to calculate $\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle F$, we only need to count all the partitions $\mu\vdash d$ such that: $\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}\leq\mu_{i}\leq\min\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}$ for all $i\geq 1$. Therefore, we have that
$\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle F$ $=\langle q^{d}\rangle\Bigg(\prod_{i=1}^{\infty}\Bigg(\sum_{j=\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}}^{\min\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}}q^{j}\Bigg)\Bigg)$
$=\langle q^{d-\sum_{i=1}^{\infty}\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}}\rangle\Bigg(\prod_{i=1}^{\infty}\Bigg(\sum_{j=\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}}^{\min\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}}q^{j-\sum_{i=1}^{\infty}\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}}\Bigg)\Bigg)$
$=\bigg\langle q^{d-\sum_{i=1}^{\infty}\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}}\bigg\rangle\Bigg(\prod_{i=1}^{\infty}\Bigg(\sum_{j=0}^{\min\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}-\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}}q^{j}\Bigg)\Bigg).$
Lemma 5.2.
Let $d,a,b,p,q\in{\mathbb{Z}}_{\geq 0}$ be nonnegative integers. Then we have
$\sum_{\mu\vdash d}\{s_{\mu}\cdot h_{a}\}_{\lambda_{1}=p}\otimes\{s_{\mu}\cdot h_{b}\}_{\lambda_{1}=q}=\sum_{\mu\vdash\big(d+a+b-\max\{p,q\}\big)}\{s_{\mu}\cdot h_{\max\{p,q\}-b}\}_{\lambda_{1}=p}\otimes\{s_{\mu}\cdot h_{\max\{p,q\}-a}\}_{\lambda_{1}=q}.$
Proof.
We denote the left-hand side (resp. right-hand side) by $F$ (resp. $G$). It suffices to show that
$\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle F=\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle G$
for any pair of partitions $\lambda^{(1)}\vdash a+d$ and $\lambda^{(2)}\vdash b+d$.
Consider three conditions:
  • $\lambda_{1}^{(1)}=p$,
  • $\lambda_{1}^{(2)}=q$, and
  • for all $i\geq 1$ we have $\min\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}\geq\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}$.
If one of them is not satisfied, Lemma 5.1 indicates that $\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle F=0=\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle G$. We henceforth suppose that all these three conditions hold. Then Lemma 5.1 indicates that
$\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle F=\Big\langle q^{d-\sum_{i=1}^{\infty}\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}}\Big\rangle f$
and
$\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle G=\Big\langle q^{d+a+b-\max\{p,q\}-\sum_{i=1}^{\infty}\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}}\Big\rangle f$
where $f=\prod_{i=1}^{\infty}\Bigg(\sum_{j=0}^{\min\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}-\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}}q^{j}\Bigg)$. Therefore, it remains to show that
$\Big\langle q^{d-\sum_{i=1}^{\infty}\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}}\Big\rangle f=\Big\langle q^{d+a+b-\max\{p,q\}-\sum_{i=1}^{\infty}\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}}\Big\rangle f.$
Note that $f$ is actually a product of finitely many palindromic polynomials in $q$, $f$ itself is also a palindromic polynomial in $q$. Thus, it suffices to show that
$\Big(d-\sum_{i=1}^{\infty}\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}\Big)+\Big(d+a+b-\max\{p,q\}-\sum_{i=1}^{\infty}\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}\Big)=\deg(f),$
which is equivalent to
(5.1) $2d+a+b=\max\{p,q\}+\sum_{i=1}^{\infty}\min\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}+\sum_{i=1}^{\infty}\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}.$
Since the right-hand side of Equation (5.1) equals
$\max\{\lambda_{1}^{(1)},\lambda_{1}^{(2)}\}+\sum_{i=1}^{\infty}\min\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}+\sum_{i=1}^{\infty}\max\{\lambda_{i+1}^{(1)},\lambda_{i+1}^{(2)}\}$
$=$ $\sum_{i=1}^{\infty}\min\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}+\sum_{i=1}^{\infty}\max\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}$
$=$ $\sum_{i=1}^{\infty}\Big(\min\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}+\max\{\lambda_{i}^{(1)},\lambda_{i}^{(2)}\}\Big)$
$=$ $\sum_{i=1}^{\infty}\Big(\lambda_{i}^{(1)}+\lambda_{i}^{(2)}\Big)=\lvert\lambda^{(1)}\rvert+\lvert\lambda^{(2)}\rvert=(a+d)+(b+d)=2d+a+b,$
which equals the left-hand side of Equation (5.1). Therefore, Equation (5.1) holds, completing our proof. ∎
Corollary 5.3.
(5.2) $\sum_{d=0}^{\min\{m,n\}}\Bigg\{\sum_{\mu\vdash d}(s_{\mu}\cdot h_{n-d})\otimes(s_{\mu}\cdot h_{m-d})\Bigg\}_{\lambda_{1}\leq n+m-d-r}=\sum_{r^{\prime}=r}^{\min\{m,n\}}\sum_{\mu\vdash r^{\prime}}(s_{\mu}\cdot h_{n-r^{\prime}})\otimes(s_{\mu}\otimes h_{m-r^{\prime}})$
Proof.
Pieri’s rule and Lemma 5.2 indicate that the left-hand side equals
$\sum_{d=0}^{\min\{m,n\}}\sum_{\mu\vdash d}\{s_{\mu}\cdot h_{n-d}\}_{n-d\leq\lambda_{1}\leq n+m-d-r}\otimes\{s_{\mu}\cdot h_{m-d}\}_{m-d\leq\lambda_{1}\leq n+m-d-r}$
$=$ $\sum_{\begin{subarray}{c}(d,p,q)\in{\mathbb{Z}}^{3}\\ 0\leq d\leq\min\{m,n\}\\ n-d\leq p\leq n+m-d-r\\ m-d\leq q\leq n+m-d-r\end{subarray}}\sum_{\mu\vdash d}\{s_{\mu}\cdot h_{n-d}\}_{\lambda_{1}=p}\otimes\{s_{\mu}\cdot h_{m-d}\}_{\lambda_{1}=q}$
$=$ $\sum_{\begin{subarray}{c}(d,p,q)\in{\mathbb{Z}}^{3}\\ 0\leq d\leq\min\{m,n\}\\ n-d\leq p\leq n+m-d-r\\ m-d\leq q\leq n+m-d-r\end{subarray}}\sum_{\mu\vdash(m+n-d-\max\{p,q\})}\{s_{\mu}\cdot h_{\max\{p,q\}-m+d}\}_{\lambda_{1}=p}\otimes\{s_{\mu}\cdot h_{\max\{p,q\}-n+d}\}_{\lambda_{1}=q}$
(5.3) $=$ $\sum_{r^{\prime}=r}^{\min\{m,n\}}\!\sum_{\begin{subarray}{c}(d,p,q)\in{\mathbb{Z}}^{3}\\ 0\leq d\leq\min\{m,n\}\\ p\geq n-d\\ q\geq m-d\\ n+m-d-\max\{p,q\}=r^{\prime}\end{subarray}}\!\sum_{\mu\vdash(m+n-d-\max\{p,q\})}\!\{s_{\mu}\cdot h_{\max\{p,q\}-m+d}\}_{\lambda_{1}=p}\otimes\{s_{\mu}\cdot h_{\max\{p,q\}-n+d}\}_{\lambda_{1}=q}$
where the second equal sign arises from Lemma 5.2, and the last equal sign arises from the following two facts:
  • $p,q\leq n+m-d-r$ if and only if $n+m-d-\max\{p,q\}\geq r$, and
  • $p\geq n-d$ and $q\geq m-d$ together indicate that $n+m-d-\max\{p,q\}\leq\min\{m,n\}$.
Note that we can eliminate $d$ in the expression (5.3) using the condition $n+m-d-\max\{p,q\}=r^{\prime}$ under the summation, i.e., we can substitute $d=n+m-\max\{p,q\}-r^{\prime}$ in the expression (5.3). Therefore, the left-hand side of Equation (5.2) equals
(5.4) $\sum_{r^{\prime}=r}^{\min\{m,n\}}\!\sum_{\begin{subarray}{c}(p,q)\in{\mathbb{Z}}^{2}\\ 0\leq n+m-\max\{p,q\}-r^{\prime}\leq\min\{m,n\}\\ p\geq\max\{p,q\}+r^{\prime}-m\\ q\geq\max\{p,q\}+r^{\prime}-n\\ \end{subarray}}\!\sum_{\mu\vdash r^{\prime}}\!\{s_{\mu}\cdot h_{n-r^{\prime}}\}_{\lambda_{1}=p}\otimes\{s_{\mu}\cdot h_{m-r^{\prime}}\}_{\lambda_{1}=q}.$
Claim: All three inequalities under the second summation in (5.4) can be discarded.
In fact, for any non-empty term $\{s_{\mu}\cdot h_{n-r^{\prime}}\}_{\lambda_{1}=p}\otimes\{s_{\mu}\cdot h_{m-r^{\prime}}\}_{\lambda_{1}=q}$ in (5.4), Pieri’s rule yields that
(5.5) $\max\{n-r^{\prime},\mu_{1}\}\leq p\leq\mu_{1}+n-r^{\prime}\,\,\,(\leq n)$
and
(5.6) $\max\{m-r^{\prime},\mu_{1}\}\leq q\leq\mu_{1}+m-r^{\prime}\,\,\,(\leq m).$
Then
(5.7) $\max\{m,n\}-r^{\prime}\leq\max\{p,q\}\leq\max\{m,n\}.$
Therefore, $0\leq\min\{m,n\}-r^{\prime}=n+m-\max\{m,n\}-r^{\prime}\leq n+m-\max\{p,q\}-r^{\prime}\leq n+m-(\max\{m,n\}-r^{\prime})-r^{\prime}=\min\{m,n\}$, so we have $0\leq n+m-\max\{p,q\}-r^{\prime}\leq\min\{m,n\}$ and hence the first equality under the summation in (5.4) can be discard. Now we suppose that $p\leq q$ without loss of generality. Then (5.6) indicates that $\max\{p,q\}+r^{\prime}-m=q+r^{\prime}-m\leq\mu_{1}$, and (5.5) indicates that $p\geq\mu_{1}$. Consequently, we have $p\geq\max\{p,q\}+r^{\prime}-m$, so we can discard the second inequality under the summation of (5.4). Additionally, $\max\{p,q\}+r^{\prime}-n=q+r^{\prime}-n\leq q$ because $r^{\prime}\leq\min\{m,n\}$. Thus, we can discard the last inequality under the summation of (5.4). Therefore, the claim above holds.
By this claim, the expression (5.4) equals the right-hand side of Equation (5.2), finishing the proof. ∎
Recall that we have studied the module structure of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}$ in Section 3.1 and note that $I_{n,m}\subseteq I_{n,m,r}$. This observation reveals an ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-equivariant surjection ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}\twoheadrightarrow{\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$. Thanks to this surjection, we finally obtain the graded ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{n}$-module structure of $R({\mathcal{UZ}}_{n,m,r})={\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$:
Theorem 5.4.
${\mathrm{grFrob}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r};q)=\sum_{d=0}^{\min\{m,n\}}q^{d}\cdot\Bigg\{\sum_{\mu\vdash d}(s_{\mu}\cdot h_{n-d})\otimes(s_{\mu}\cdot h_{m-d})\Bigg\}_{\lambda_{1}\leq n+m-d-r}$
Proof.
The containment $I_{n,m}\subseteq I_{n,m,r}$ indicates an ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-equivariant surjection
${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m}\twoheadrightarrow{\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r},$
which, together with Proposition 3.5, means that
${\mathrm{grFrob}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r};q)\leq{\mathrm{grFrob}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}])/I_{n,m};q)=\sum_{d=0}^{\min\{m,n\}}q^{d}\cdot\sum_{\mu\vdash d}(s_{\mu}\cdot h_{n-d})\otimes(s_{\mu}\cdot h_{m-d}).$
Then Lemma 2.2 and Corollary 3.9 indicates that
(5.8) ${\mathrm{grFrob}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r};q)\leq\sum_{d=0}^{\min\{m,n\}}q^{d}\cdot\Bigg\{\sum_{\mu\vdash d}(s_{\mu}\cdot h_{n-d})\otimes(s_{\mu}\cdot h_{m-d})\Bigg\}_{\lambda_{1}\leq n+m-d-r}.$
However, Inequality (5.8) indicates that
(5.9) ${\mathrm{Frob}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r})$ $\leq\sum_{d=0}^{\min\{m,n\}}\Bigg\{\sum_{\mu\vdash d}(s_{\mu}\cdot h_{n-d})\otimes(s_{\mu}\cdot h_{m-d})\Bigg\}_{\lambda_{1}\leq n+m-d-r}$
$=\sum_{r^{\prime}=r}^{\min\{m,n\}}\sum_{\mu\vdash r^{\prime}}(s_{\mu}\cdot h_{n-r^{\prime}})\otimes(s_{\mu}\otimes h_{m-r^{\prime}})$
$={\mathrm{Frob}}\Bigg(\bigoplus_{r^{\prime}=r}^{\min\{m,n\}}{\mathbb{C}}[{\mathcal{Z}}_{n,m,r^{\prime}}]\Bigg)={\mathrm{Frob}}({\mathbb{C}}[{\mathcal{UZ}}_{n,m,r}])$
where the first equal sign derives from Corollary 5.3, and the second equal sign arises from Equation (3.3). Now Theorem 4.12 tells us $R({\mathcal{UZ}}_{n,m,r})={\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$, so orbit harmonics yields an ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-module isomorphism
${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}\cong{\mathbb{C}}[{\mathcal{UZ}}_{n,m,r}],$
forcing all the inequalities in (5.9) to be equalities, and hence forcing Inequality (5.8) to be the equality that we desire. ∎
6. Conclusion
We provide an open problem regarding log-concavity. Recall that a sequence of positive real numbers $(a_{1},a_{2},\cdots,a_{N})$ is log-concave if for all $1<i<N$ we have $a_{i}^{2}\geq a_{i-1}\cdot a_{i+1}$. Chen conjectured [1] that the sequence $(a_{n,1},a_{n,2},\cdots,a_{n,n})$ is log-concave where $a_{n,k}$ is the number of permutations in ${\mathfrak{S}}_{n}$ with longest subsequence of length $k$. Using the terminology of orbit harmonics, Chen’s conjecture means that the Hilbert series of the orbit harmonics ring $R({\mathfrak{S}}_{n})$ of the permutation matrix locus ${\mathfrak{S}}_{n}$ is log-concave.
We can generalize the definition of log-concavity by incorporating group action. Let $G$ be a group and $(V_{1},V_{2},\cdots,V_{N})$ is a sequence of $G$-modules. We say that $(V_{1},V_{2},\cdots,V_{N})$ is $G$-equivariant log-concave if for all $1<i<N$ we have a $G$-module surjection $V_{i}\otimes V_{i}\twoheadrightarrow V_{i-1}\otimes V_{i+1}$ where the $G$-action is diagonal, i.e., $g\cdot(v\otimes w)=(g\cdot v)\otimes(g\cdot w)$. A graded $G$-module $V=\bigoplus_{d=0}^{N}V_{d}$ is $G$-equivariant log-concave if the sequence $(V_{1},V_{2},\cdots,V_{N})$ is $G$-equivariant log-concave. Rhoades conjectured [13] that $R({\mathfrak{S}}_{n})$ is ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{n}$-equivariant log-concave, which generalizes Chen’s conjecture.
Our locus ${\mathcal{UZ}}_{n,m,r}$ coincides with the permutation matrix locus ${\mathfrak{S}}_{n}$ visited by Rhoades if $n=m=r$. Therefore, we raise the following conjecture, which further generalizes Rhoades’s conjecture.
Conjecture 6.1.
The graded ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-module $R({\mathcal{UZ}}_{n,m,r})$ is ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-log-concave.
Conjecture 6.1 has been checked by coding for $n\leq 8,m\leq 10$. The more flexibility provided by Conjecture 6.1 than Rhoades’s conjecture, such as three parameters $n,m,r$ rather than $n$ itself, may permit more potential induction strategies and make the proof of log-concavity easier.
Another interesting future direction is to study the orbit harmonics ring $R({\mathcal{Z}}_{n,m,r})$. Since ${\mathcal{Z}}_{n,m,r}\subseteq{\mathcal{UZ}}_{n,m,r}$, we need to add some elements to the generating set in Definition 3.1 to obtain the defining ideal ${\mathrm{gr}}{\mathbf{I}}({\mathcal{Z}}_{n,m,r})$, such as $\sum_{i=1}^{n}\sum_{j=1}^{m}x_{i,j}$ and ${\mathfrak{m}}({\mathcal{R}})$ for ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r+1}$. Some relevant results will be presented in a forthcoming paper.
Finally, we may also consider extending the results to colored rook placement loci. Let $N,k$ be positive integers, the $k$-colored permutation group ${\mathfrak{S}}_{N,k}$ is the wreath product $({\mathbb{Z}}/r{\mathbb{Z}})\wr{\mathfrak{S}}_{N}$ which can be embedded into ${\mathrm{Mat}}_{N\times N}({\mathbb{C}})$ by
[…]
Define the $k$-colored rook placement loci ${\mathcal{Z}}_{n,m,r}^{(k)}$ by
[…]
and let
${\mathcal{UZ}}_{n,m,r}^{(k)}\coloneqq\bigsqcup_{r^{\prime}=r}^{\min\{m,n\}}{\mathcal{Z}}_{n,m,r^{\prime}}^{(k)}.$
Problem 6.2.
Find the graded ${\mathfrak{S}}_{n,k}\times{\mathfrak{S}}_{m,k}$-module structures of $R({\mathcal{Z}}_{n,m,r}^{(k)})$ and $R({\mathcal{UZ}}_{n,m,r}^{(k)})$.
7. Acknowledgements
The authors are thankful to Brendon Rhoades for constructive suggestions about the content and structure of this paper.